OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6).

OpenStudy (anonymous):

y = x + 7 y = −3x + 3 y = x + 3 y = −3x + 7 these are the answer choices

OpenStudy (souvik):

what is the slope of this equation?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):



OpenStudy (anonymous):